OpenStudy (anonymous):

(5x-3)^2 = 36 using the square root property

OpenStudy (anonymous):

\[6x-3 = \pm6\] x1 = 9/5 x2 = -3/5

OpenStudy (anonymous):

Sorry, i meant 5x-3, not 6x. In the solution it's correct (divided by 5).